OpenStudy (anonymous):

Simplify the following expression: (x + 6y) - (3x − 10y). If the final answer is written in the form Ax + By, what is the value of A? Numerical Answers Expected! Answer for Blank 1: I have literally no idea what to do...

OpenStudy (aravindg):

simple just add like terms

OpenStudy (aravindg):

x-3x+6y+10y= ?

OpenStudy (anonymous):

so its just the same as if they gave me choices i'm just writing my answer?

OpenStudy (aravindg):

what did you get ?

OpenStudy (anonymous):

2x-4y

OpenStudy (aravindg):

try again ? x-3x+6y+10y= ?

OpenStudy (aravindg):

(1-3)x+(6+10)y = ?

OpenStudy (anonymous):

-2x+16y?

OpenStudy (aravindg):

\[\huge \color{red}{\checkmark }\]

OpenStudy (anonymous):

yay! thank you!

OpenStudy (aravindg):

yw :)